OpenStudy (anonymous):

The sum of two numbers is 55. One number is 5 more than the other. What is the smaller number?

OpenStudy (apoorvk):

let one no. be x. therefore other is x+5. then (x) + (x+5)=55 solve and enjoy!

OpenStudy (eyust707):

heres how he got that: the sum of two numbers is 55; \[x+y=55\] one is 5 more than the other; \[x+5=y\] substitute in for y \[x+(x+5)=55\] then solve

OpenStudy (anonymous):

thanks!!!